OpenStudy (anonymous):

solve using multiplication principle.7s=42/19

OpenStudy (saifoo.khan):

jt, the answer for last one is: n = 12/5

OpenStudy (anonymous):

k thanks

OpenStudy (saifoo.khan):

\[s = \frac{42}{19} \times \frac17\]Multiply them horizontally.